# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ Appendix B Analysis of Feature Reconstruction
Let $r_{i}=\frac{2}{K}-(y_{i1}+y_{i2})$, then $r_{i}=\frac{2}{K}$ or $\frac{2-K}{K}$, which is non-zero. Then $g({\bm{w}}_{j})=\frac{1}{m}\sum_{i=1}^{B}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})$. We can use the same attacking method in (Wang et al., 2023) that conduct noisy tensor decomposition to $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{3}({\bm{w}}_{j})$, where $H_{p}$ is the $p$-th Hermite function. In our method, we will mainly use $H_{2}({\bm{w}})={\bm{w}}{\bm{w}}^{\top}-I$ and $H_{3}({\bm{w}})={\bm{w}}^{\otimes 3}-{\bm{w}}\tilde{\otimes}I$, where ${\bm{w}}\tilde{\otimes}I(i,j,k)=w_{i}\delta_{jk}+w_{j}\delta{ki}+w_{k}\delta{ij}$.
Theorem B.1.
For a 2-layer neural network with input dimension $d$, width $m$, and cross-entropy loss with $K$ classes, under Assumption A.1, we observe gradient descent steps with batch size $K^{2}B^{4}\leq\tilde{O}(d)$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound:
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
with high probability.
Proof.
We choose weights as discussed above. Let $\hat{P}=\sum_{j=1}^{m}g({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\hat{{\bm{T}}}=\sum_{j=1}^{m}g({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. We also denote $P=m\mathbb{E}_{{\bm{w}}\sim\mathcal{N}(0,I)}g({\bm{w}})H_{2}({\bm{w}})$ and ${\bm{T}}=m\mathbb{E}_{{\bm{w}}\sim\mathcal{N}(0,I)}g({\bm{w}})H_{3}({\bm{w}})$. Note that $H_{2}({\bm{w}})={\bm{w}}{\bm{w}}^{\top}-I$ and $H_{3}({\bm{w}})={\bm{w}}^{\otimes 3}-{\bm{w}}\tilde{\otimes}I$. Then by Lemma 3.1, we have $P=\sum_{i=1}^{B}(\frac{2}{K}-y_{1i}-y_{2i})\nu{\bm{x}}_{i}^{\otimes 2}$ and ${\bm{T}}=\sum_{i=1}^{B}(\frac{2}{K}-y_{1i}-y_{2i})\lambda{\bm{x}}_{i}^{\otimes 3}$. In the following proof, we give the bound of $\|\hat{P}-P\|$ and $\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|$.
Error bound of $\hat{P}$. Denote $Z_{j}=\sigma({\bm{w}}_{j}^{\top}{\bm{x}})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$. We check the condition of Lemma A.2. For any ${\bm{x}}$
(I) We have the bound of the norm of $Z_{j}$:
$\left\|Z_{j}\right\|\leq\left|\sigma({\bm{w}}_{j}^{\top}{\bm{x}})\right|(\|{\bm{w}}_{j}\|^{2}+1)\leq Cd\log(8Bm/\delta)$
with probability $1-\frac{\delta}{4Bm}$.
(II) We have
$\max\left\{\left\|\mathbb{E}\left[Z^{\top}Z\right]\right\|,\left\|\mathbb{E}\left[ZZ^{\top}\right]\right\|\right\}=\left\|\mathbb{E}\left[Z^{2}\right]\right\|=\left\|\mathbb{E}_{{\bm{w}}\sim\mathcal{N}(0,I)}\sigma({\bm{w}}^{\top}{\bm{x}})^{2}({\bm{w}}{\bm{w}}^{T}-I)^{2}\right\|.$
By the symmetry of normal distribution, we can assume ${\bm{x}}={\bm{e}}_{1}$. Then we have to bound $\|Q\|:=\left\|\mathbb{E}\sigma(w_{1})^{2}({\bm{w}}{\bm{w}}^{\top}-I)^{2}\right\|$, where we can directly compute that $Q$ is a diagonal matrix with positive elements no larger than $(d+1)\max\left\{\mathbb{E}\sigma(w_{1})^{2}w_{1}^{2},\mathbb{E}\sigma(w_{1})^{2}\right\}$. Since $\sigma$ is 1-Lipschitz, $\|B\|\leq O(d)$.
$\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}=\mathbb{E}\sigma({\bm{w}}^{\top}{\bm{x}})^{2}({\bm{a}}^{\top}({\bm{w}}{\bm{w}}^{\top}-I){\bm{a}})^{2}.$
Similar to (II), we can assume that ${\bm{x}}={\bm{e}}_{1}$. Then we directly compute the expectation:
$\begin{split}\mathbb{E}({\bm{a}}^{\top}Z{\bm{a}})^{2}&=\mathbb{E}\sigma(w_{1})^{2}\left(\sum_{i=1}^{d}a_{i}^{2}(w_{i}^{2}-1)+\sum_{i\neq j}a_{i}a_{j}w_{i}w_{j}\right)^{2}\\ &=\mathbb{E}\left[\sum_{i=1}^{d}\sigma(w_{1})^{2}a_{i}^{4}(w_{i}^{2}-1)^{2}+\sum_{i\neq j}\sigma(w_{1})^{2}a_{i}^{2}a_{j}^{2}w_{i}^{2}w_{j}^{2}\right]\\ &\lesssim\left(\sum_{i=1}^{d}a_{i}^{2}\right)^{2}=1.\end{split}$ (9)
Thus, $\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}(\mathbb{E}({\bm{a}}^{\top}Z\bm{b})^{2})^{1/2}\leq O(1)$.
Moreover, we have to bound $\|\mathbb{E}[Z]\|$, and we have
$\mathbb{E}\left[Z\right]=\mathbb{E}\sigma({\bm{w}}^{\top}{\bm{x}})({\bm{w}}{\bm{w}}^{\top}-I)=\mathbb{E}\sigma^{\prime\prime}({\bm{w}}^{\top}{\bm{x}}){\bm{x}}{\bm{x}}^{\top}=\mathbb{E}_{z\sim\mathcal{N}(0,1)}\sigma^{\prime\prime}(z){\bm{x}}{\bm{x}}^{\top}$
by Lemma 3.1. Then $\|\mathbb{E}[Z]\|\lesssim\|{\bm{x}}{\bm{x}}^{\top}\|\leq O(1)$.
Thus, by Theorem A.2, for any $i$ we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)-\mathbb{E}r_{i}\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|\lesssim|r_{i}|\sqrt{\frac{d\log(4B/\delta)\log(8Bm/\delta)}{m}}$
with probability $1-\frac{\delta}{2B}$. Then with probability $1-\frac{\delta}{2}$, we have
$\begin{split}\|P-\hat{P}\|&\leq\sum_{i=1}^{B}\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)-\mathbb{E}r_{i}\sigma({\bm{w}}^{\top}{\bm{x}}_{i})({\bm{w}}{\bm{w}}^{\top}-I)\right\|\\ &\lesssim\sqrt{\frac{d\log(4B/\delta)\log(16Bm/\delta)}{m}}\sum_{i=1}^{B}|r_{i}|\leq\tilde{O}(\frac{B\sqrt{d}}{\sqrt{m}}).\end{split}$ (10)
Error bound of $\hat{{\bm{T}}}(V,V,V)$. Let $R\in\mathbb{R}^{B\times B^{2}}$ is the flatten along the first dimension of ${\bm{T}}(V,V,V)$. Since for any symmetric third order tensor ${\bm{E}}$ and its flatten along the first dimension $E^{(1)}$ we have $\|{\bm{E}}\|\leq\|E^{(1)}\|$, we can bound $\|R-\hat{R}\|$ instead.
Denote ${\bm{T}}_{i}=\frac{1}{m}\sum_{j=1}^{m}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{i}\tilde{\otimes}I)$ and ${R_{i}}=T_{i}^{(1)}(V,V,V)$. For any $i$, let $\tilde{{\bm{Z}}}_{ij}=\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{i}\tilde{\otimes}I)$ and $Z_{ij}=\tilde{{\bm{Z}}}^{(1)}_{ij}(V,V,V)$, then $R_{i}=\frac{1}{m}\sum_{j=1}^{m}Z_{ij}$. We check the conditions of Theorem A.2. For any ${\bm{x}}$
(I) We have
$\left\|Z_{j}\right\|\lesssim|\sigma({\bm{w}}_{j}^{\top}{\bm{x}})|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{B^{3/2}}{\sqrt{m}})$
with probability $1-\frac{\delta}{4Bm}$.
(II) We have
$\max\left\{\left\|\mathbb{E}\left[Z^{\top}Z\right]\right\|,\left\|\mathbb{E}\left[ZZ^{\top}\right]\right\|\right\}\leq\mathbb{E}\left[\left\|Z\right\|^{2}\right]\lesssim\mathbb{E}\left[(\sigma({\bm{w}}_{j}^{\top}{\bm{x}}))^{4}\right]^{1/2}\mathbb{E}\left[\|V^{\top}{\bm{w}}_{j}\|^{12}\right]^{1/2}\lesssim B^{3}.$
(III) We have
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}\left({\mathbb{E}}\left[\left({\bm{a}}^{\top}Z\bm{b}\right)^{2}\right]\right)^{1/2}\leq\left(\mathbb{E}\left[\left\|Z\right\|^{2}\right]\right)^{1/2}\lesssim B^{3/2}.$
Additionally, we also have to bound $\left\|\mathbb{E}\left[Z\right]\right\|$.
$\begin{split}\mathbb{E}[Z]&=\sum_{i=1}^{B}r_{i}^{*}\left(\mathbb{E}[\sigma^{(3)}({\bm{w}}^{\top}{\bm{x}}){\bm{x}}^{\otimes 3}]\right)^{(1)}(V,V,V)\\ &=\sum_{i=1}^{B}\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma^{(3)}(z)](V^{\top}{\bm{x}})\mathrm{vec}\left[(V^{\top}{\bm{x}})(V^{\top}{\bm{x}})^{\top}\right]^{\top}\end{split}$ (11)
by Lemma 3.1. Then by Theorem A.2, we have for any $i$ that
$\|R_{i}-\mathbb{E}R_{i}\|\leq|r_{i}|\tilde{O}(\frac{B^{3/2}}{\sqrt{m}})$
with probability $1-\frac{\delta}{2B}$. Then with probability $1-\frac{\delta}{2}$, we have
$\|{\bm{T}}(V,V,V)-\hat{{\bm{T}}}(V,V,V)\|\leq\sum_{i=1}^{B}\|R_{i}-\mathbb{E}R_{i}\|\leq\tilde{O}(\frac{B^{5/2}}{\sqrt{m}}).$ (12)
Since $\frac{1}{K}\leq|r_{i}|\leq 1$, we have that $P$’s smallest component $\nu_{\mathrm{min}}\geq\frac{|\nu|}{K}$, ${\bm{T}}$’s smallest component $\lambda_{\mathrm{min}}\geq\frac{|\lambda|}{K}$ and $\kappa\leq K$. Then by Theorem A.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}_{i}}\|\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
for all $i=1,\dots,B$ with high probability. ∎
Remark B.2.
When $K>3$ is odd, we can similarly set $\|U\|=\frac{K+1}{2}$ with all other settings not changed. Then $p_{1}+p_{2}=\frac{2(e^{b_{1}z}+e^{b_{2}z})}{(K+1)e^{b_{1}z}+(K-1)e^{b_{2}z}}$ and $\frac{2}{K+1}\leq p_{1}+p_{2}\leq\frac{2}{K-1}$. Thus, $|r_{i}|\geq\min\{\frac{2}{K+1},\frac{3-K}{K-1}\}\geq\frac{1}{K}$.
Remark B.3.
For $K=2,3$, we can set $c_{k}=c$ for all $k=1,\dots,K$, then $p_{k}=\frac{1}{K}$ for all $k$. Then we construct $g({\bm{w}}_{j})$ with the first class:
$g({\bm{w}}_{j})=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{1j}}=\frac{1}{m}\sum_{i=1}^{B}(\frac{1}{K}-y_{i1})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i}).$
We still have $|r_{i}|\geq\frac{1}{K}$ with this construction. In fact, this method works for any $K\geq 2$. However, setting all the weights $c_{k}$ to be the same will cause
$\frac{\partial\ell}{\partial{\bm{w}}_{j}}=\sum_{k=1}^{K}\frac{\partial\ell}{\partial z_{k}}\frac{\partial z_{k}}{\partial{\bm{w}}_{j}}=\frac{ca_{j}}{m}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}){\bm{x}}\cdot\sum_{k=1}^{K}(p_{k}-y_{k})=0.$
Then one cannot reconstruct training samples with the gradient of $W$. Thus, we will apply the definition in (8) in the future discussion.
Appendix C Proofs of Defending Strategies
@@ C.1 Local Aggregation
For local aggregation, the observation is $\Theta^{(2)}-\Theta^{(0)}$. In our attack method, we conduct tensor decomposition to $\tilde{g}=-\sum_{j=1}((a_{1j}^{(2)}-a_{1j}^{(0)})+(a_{2j}^{(2)}-a_{2j}^{(0)}))H_{3}({\bm{w}}_{j})$, where $H_{p}$ is the $p$th Hermite function. Then we have the following reconstruction error bounds under the setting of two gradient descent steps with a same batch of input and with two different batches input.
Proposition C.1.
For a 2-layer neural network with input dimension $d$, width $m$ and cross-entropy loss with $K$ classes, under Assumption A.1, the observed update $\Theta^{(2)}-\Theta^{(0)}$ is the result of 2 gradient descent steps trained a same batch with size $B$, where $K^{2}B^{4}\leq\tilde{O}(d)$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$ and learning rate $\eta=O(1)$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound:
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$ (13)
@@
For a 2-layer neural network with input dimension $d$, width $m$ and cross-entropy loss with $K$ classes, under Assumption A.1, the observed update $\Theta^{(2)}-\Theta^{(0)}$ is the result of 2 gradient descent steps, where the first step is trained with $B$ data and the second step is trained with $N-B$ data. Here $B<N\leq 2B$ and $K^{2}B^{4}\leq\tilde{O}(d)$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$ and learning rate $\eta=O(1)$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound:
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(N\sqrt{\frac{d}{m}})$ (14)
@@
The first set of lemmas show that ${\bm{w}}^{(1)}$ has a similar effect to ${\bm{w}}$ in the reconstruction.
@@
If the learning rate $\eta=O(1)$, we have
$|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\lesssim\frac{BK}{m}$ (15)
for any $i$ and $j$.
@@
Note that ${\bm{w}}_{j}^{(1)}={\bm{w}}_{j}-\frac{\eta}{m}\sum_{i=1}^{B}\sum_{k=1}^{K}(p_{ki}-y_{ki})a_{kj}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}$. Then
$\begin{split}&|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\\ =&\left|\sigma\left({\bm{w}}_{j}^{\top}{\bm{x}}_{i}-\frac{\eta}{m}\sum_{i=1}^{B}\sum_{k=1}^{K}(p_{ki}-y_{ki})a_{kj}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right)-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|\\ \leq&\frac{\eta}{m}\sum_{i=1}^{B}\sum_{k=1}^{K}(p_{ki}-y_{ki})a_{kj}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\lesssim\frac{BK}{m}\end{split}$ (16)
for all $i,\ j$. ∎
@@
If the learning rate $\eta=O(1)$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}\left[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})\right]\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\leq\tilde{O}(\frac{BKd}{m})$ (17)
@@
We define $Z_{j}=[(p^{(1)}_{1i}-y_{1i})+(p^{(1)}_{2i}-y_{2i})]|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, then we have

### v2

@@ B.4 Reconstruction Upper Bounds
For each ${\bm{w}}_{j}$, we define $g({\bm{w}}_{j})=\sum_{i=1}^{B}r_{i}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})$. We can use the same attacking method in [Wang et al., 2023] that conduct noisy tensor decomposition to $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{3}({\bm{w}}_{j})$, where $H_{p}$ is the $p$-th Hermite function. In our method, we will mainly use $H_{2}({\bm{w}})={\bm{w}}{\bm{w}}^{\top}-I$ and $H_{3}({\bm{w}})={\bm{w}}^{\otimes 3}-{\bm{w}}\tilde{\otimes}I$, where ${\bm{w}}\tilde{\otimes}I(i,j,k)=w_{i}\delta_{jk}+w_{j}\delta{ki}+w_{k}\delta{ij}$.
Lemma B.3 (Stein’s Lemma).
$\mathbb{E}[g(X)H_{p}(X)]=\mathbb{E}[g^{(p)}(X)],$ (5)
We define
$\hat{P}=\frac{1}{m}\sum_{j=1}^{m}g({\bm{w}}_{j})H_{2}({\bm{w}}_{j})$ (6)
$P=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{B}r_{i}^{*}\sigma^{\prime\prime}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}{\bm{x}}_{i}^{\top}\right]$ (7)
$\hat{{\bm{T}}}=\frac{1}{m}\sum_{j=1}^{m}g({\bm{w}}_{j})H_{3}({\bm{w}}_{j})$ (8)
${\bm{T}}=\operatorname{\mathbb{E}}\left[\sum_{i=1}^{B}r_{i}^{*}\sigma^{(3)}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}^{\otimes 3}\right].$ (9)
Then with Lemma B.3 and concentration bounds, we have the following lemmas.
Lemma B.4 (Proposition 5.5 in [Wang et al., 2023]).
Under Assumption B.1 and B.2, $|y_{i}|\leq 1$, then for $\delta\leq\frac{2}{d}$ and $m\gtrsim\log(8/\delta)$, we have
$\|\hat{P}-P\|\leq\tilde{O}(\frac{B\sqrt{d}}{\sqrt{m}})$
with probability $1-\delta$.
Lemma B.5 (Proposition 5.6 in [Wang et al., 2023]).
Under Assumption B.1 and B.2, $|y_{i}|\leq 1$, and $\|VV^{\top}-UU^{\top}\|\leq 1/4$, then for $\delta\leq\frac{2}{B}$ and $m\gtrsim\log(6/\delta)$ we have
$\|\bar{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\tilde{O}(\frac{B^{5/2}}{\sqrt{m}})$
with probability $1-\delta$.
Combining with Theorem B.1, we have a upper bound for data reconstruction error.
Theorem B.6 (Theorem 5.1 in [Wang et al., 2023], full version of Theorem 3.1).
Under Assumption B.1 and B.2, $y_{i}\in\{\pm 1\}$, if we have $B\leq\tilde{O}(d^{1/4})$ and $m\geq\tilde{\Omega}(\frac{d}{\min\{\nu^{2},\lambda^{2}\}\pi_{\min}^{4}})$, then for $\delta\leq\frac{2}{d}$ and $m\gtrsim\log(8/\delta)$, the output of tensor based reconstruction satisfies:
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(\sqrt{\frac{d}{m}})$
with probability $1-\delta$.
Appendix C Proofs of Upper Bounds with Defenses
@@ C.1 Local Aggregation
For local aggregation, the observation is $\Theta^{(2)}-\Theta^{(0)}$. In our attack method, we conduct tensor decomposition to $\tilde{g}=-\sum_{j=1}^{m}(a_{j}^{(2)}-a_{j}^{(0)})H_{3}({\bm{w}}_{j})$. Here a notation with superscript $(k)$ means the corresponding value after $k$ steps of iteration. Then we have the following reconstruction error bounds under the setting of two gradient descent steps with a same batch of input and with two different batches input.
Proposition C.1 (Full version of Proposition 3.1).
Under Assumption B.1 and B.2, $y_{i}\in\{\pm 1\}$, the observed update $\Theta^{(2)}-\Theta^{(0)}$ is the result of 2 gradient descent steps trained a same batch with size $B$, where $B^{4}\leq\tilde{O}(d)$. If $m\geq\tilde{\Omega}(\frac{B^{2}d}{\nu^{2}\pi_{\min}^{2}})$ and we set learning rate for two layers with different scales $\eta_{a}=O(\frac{1}{m^{2}})$ and $\eta_{w}=O(1)$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound:
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$ (10)
@@
Under Assumption B.1 and B.2, $y_{i}\in\{\pm 1\}$, the observed update $\Theta^{(2)}-\Theta^{(0)}$ is the result of 2 gradient descent steps, where the first step is trained with $B$ data and the second step is trained with $N-B$ data. Here $B<N\leq 2B$ and $B^{4}\leq\tilde{O}(d)$. If $m\geq\tilde{\Omega}(\frac{B^{2}d}{\nu^{2}\pi_{\min}^{2}})$ and we set learning rate for two layers with different scales $\eta_{a}=O(\frac{1}{m^{2}})$ and $\eta_{w}=O(1)$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound:
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(N\sqrt{\frac{d}{m}})$ (11)
@@
Remark C.3.
In these propositions we assume that the learning rate for the two layers are different since $a_{j}\sim\mathcal{N}(0,\frac{1}{m^{2}})$ but ${\bm{w}}_{j}\sim\mathcal{N}(0,I_{d})$ have different scaling. If we use a single learning rate $\eta$ for all parameters, the results still hold if we assume $\eta=O(\frac{1}{m^{2}})$.
@@
The first lemma show that ${\bm{w}}^{(1)}$ has a similar effect to ${\bm{w}}$ in the reconstruction.
@@
If the learning rate ${\bm{w}}$ $\eta_{w}=O(1)$, we have
$|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\leq\tilde{O}(\frac{B}{m})$ (12)
for any $i$ and $j$ with probability $1-\frac{\delta}{Bm}$.
@@
Note that ${\bm{w}}_{j}^{(1)}={\bm{w}}_{j}-{\eta_{w}}\sum_{i=1}^{B}a_{j}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}$. Then
$\begin{split}&|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|\\ =&\left|\sigma\left({\bm{w}}_{j}^{\top}{\bm{x}}_{i}-\eta_{w}\sum_{i=1}^{B}a_{j}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right)-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|\\ \leq&\eta_{w}\sum_{i=1}^{B}a_{j}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\leq\tilde{O}(\frac{B}{m})\end{split}$ (13)
with probability $1-\frac{\delta}{Bm}$ for all $i,\ j$. ∎
Then we give lemmas showing that $r_{i}^{(1)}$ is very close to $r_{i}$ and the difference only changes a little in the reconstruction.
@@
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, we have
$\left|r_{i}^{(1)}-r_{i}\right|\leq\tilde{O}(\frac{B}{m})$ (14)
with probability $1-\frac{\delta}{Bm}$ for any $i$.
Proof.
We have $r_{i}=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-y_{i}$ and $r_{i}^{(1)}=\sum_{j=1}^{m}a_{j}^{(1)}\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-y_{i}$. Note that $a_{kj}^{(1)}=a_{kj}-\eta_{a}\sum_{i=1}^{B}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})$. Then by Lemma C.4 we have
$\begin{split}\left|r_{i}^{(1)}-r_{i}\right|&=\left|\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\sum_{j=1}^{m}\left(a_{j}-{\eta_{a}}\sum_{i=1}^{B}\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right)\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\right|\\ &\leq\sum_{j=1}^{m}a_{j}\left|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})\right|+\eta_{a}\sum_{j=1}^{m}\sum_{i=1}^{B}(\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i}))^{2}\\ &\quad\quad\quad\quad+{\eta_{a}}\sum_{j=1}^{m}\sum_{i=1}^{B}|\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})||\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})|\\ &\leq\tilde{O}(\frac{B}{m})+\tilde{O}(\frac{B\log(2Bm^{2}/\delta)}{m})+\tilde{O}(\frac{B\log(2Bm^{2}/\delta)}{m^{2}})\leq\tilde{O}(\frac{B}{m})\end{split}$ (15)
with probability $1-\frac{\delta}{Bm}$. ∎
Corollary C.6.
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, we have $r_{i}^{(1)}=\tilde{O}(1)$ with probability $1-\frac{\delta}{Bm}$.
Proof.
We have $r_{i}=\tilde{O}(1)$ with probability $1-\frac{\delta}{Bm}$. Then the result holds directly by Lemma C.5. ∎
Lemma C.7.
If the learning rate $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{2}})$, we have
$\left\|\frac{1}{m}\sum_{j=1}^{m}r_{i}^{(1)}\left|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\right|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\leq\tilde{O}(\frac{Bd}{m})$ (16)
@@
We define $Z_{j}=r_{i}^{(1)}|\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i})-\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$, then we have